# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

### v1

@@ Abstract
We present CD-NGP, which is a fast and scalable representation for 3D reconstruction and novel view synthesis in dynamic scenes. Inspired by continual learning, our method first segments input videos into multiple chunks, followed by training the model chunk by chunk, and finally, fuses features of the first branch and subsequent branches. Experiments on the prevailing DyNeRF dataset demonstrate that our proposed novel representation reaches a great balance between memory consumption, model size, training speed, and rendering quality. Specifically, our method consumes $85\%$ less training memory ($<14$GB) than offline methods and requires significantly lower streaming bandwidth ($<0.4$MB/frame) than other online alternatives. Furthermore, we conduct experiments on a self-collected extra-long video sequences dataset, which demonstrates the superior scalability of our method over previous state-of-the-art approaches.
@@
Continual learning Dynamic novel view synthesis Neural Radiance Field.
1 Introduction
@@
To address the dynamic reconstruction problem, amount of efforts have been made under the offline setting, requiring access to the entire dataset in the whole training stage. For example, [36, 28, 34, 33, 15, 29, 44] propose to use MLP-based representation and achieve high-quality results. [40, 17, 7, 16, 45] speed up the optimization by using tensor factorization and voxel representation, respectively. The more recent continual-learning-based methods [27, 48] can train models frame by frame, and streamable representations [43] enable the model to be loaded on the fly at inference. However, the offline methods require huge host memory e.g., taking more than $100GB$ when reconstructing multi-view dynamic videos of 300 frames, and the continual learning-based methods are weak in balancing the convergence speed and model compactness. As a result, these methods are limited in scalability, making them unsuitable for long videos.
@@
In this paper, we propose Continual Dynamic neural graphics primitives (CD-NGP), a novel, fast, and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos. As illustrated in Fig. 1, we evenly split the dynamic videos into multiple chunks based on the timestamp and conduct reconstruction with different model branches, where each branch includes MLP regressors, spatial encoding for $(x,y,z)$, and temporal encoding for $t$. We encode the different chunks by hash tables with different complexities to ensure model scalability and then continually train the model chunk by chunk to alleviate memory occupation. For each individual chunk, we locate the corresponding model branch, sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids [31], and represent spatial features by spatial hash grids. The spatial features of the first branch and the current branch are fused by element-wise sum as default or concatenation if applicable. The temporal features are encoded by the hash table for time stamp $t$ in the current branch only. Spatial and temporal features are concatenated along with viewing direction $d$ and input to MLP regressors for rendering and optimization.
@@
Extensive experiments demonstrate the great scalability of our method: compared with prevailing offline baselines, our method consumes $85\%$ less host memory at the training phase, $40\%$ less model size at the inference phase, and delivers comparable results. Compared with online baselines, our method achieves compact model size, high quality, and fast convergence simultaneously. The main contribution of this work is summarized as 3 folds:
Figure 1: The demonstration of our continual learning pipeline for dynamic scenes. Please note the dimensions of representations for different branches remain the same while the complexities of hash tables are flexible and could be configured to ensure scalability. The spatial hash features could be encoded along voxels [31], planes [17] or both [38].
@@
  • We propose a novel method combining spatial and temporal hash encodings for dynamic scenes in short intervals with high scalability.
@@
  • We propose a fast continual learning framework for dynamic scenes, producing comparable results to offline state-of-the-art methods while reducing memory and storage occupation significantly.
@@
  • We provide a dataset of extra-long video sequences ($>3$ minutes) and further demonstrate the scalability of our method on long multi-view videos.
@@ 2.1 Fast novel view synthesis in static scenes
The acceleration of NeRFs falls into two categories: training acceleration and rendering acceleration. Plenoxels [18] uses voxel grids to represent the scene and achieves magnitudes of acceleration over MLP backbones. TensoRF [9] decomposes the 3D scene as the product of tensors and achieves fast novel view synthesis. Plane-based methods [17, 21] project 3D points onto 3 orthogonal 2D planes to ensure sparsity and achieve fast convergence. Instant-NGP [31] leverages an occupancy grid and a multi-resolution hash encoding along with optimized tiny MLPs to enable fast novel view synthesis within a minute. The recent 3D Gaussian method [22] represents the scenes as a union of explicit Gaussian point clouds and achieves high fidelity, fast convergence, and real-time rendering on high-end hardware by eliminating the querying burden of spatial representations. KiloNeRF [37] distills the large MLP in NeRF [30] into tiny MLPs to accelerate rendering by thousands of times. Hedman et al. [20] propose the baking method to accelerate the rendering process. They alpha-compose the cached features along the ray and use MLPs only once per ray to overcome the computation load of the per-point MLP querying in NeRF. BakedSDF [56] stores the view-dependent features into Gaussian spheres along the reconstructed surfaces, and achieves real-time rendering on ordinary laptops. MobileNeRF[10] represents scenes into meshes and enables real-time rendering on mobile devices. Light field methods [42, 8, 19] directly represent the scene as a set of rays to eliminate volume rendering and enable real-time applications on mobile devices.
@@ 2.2 Representations of dynamic scenes
There are mainly 3 variants of NeRF for dynamic scenes: MLP, tensor factorization, and voxel representations. D-NeRF [36] leverages a large deformation MLP to model the transition and the trajectories of the objects and extends NeRF to simulated dynamic scenes. DyNeRF [28] uses long latent vectors to encode temporal information along with large MLPs to achieve high-fidelity novel view synthesis in real dynamic scenes. Park et al. [34] propose an interpolation-based MLP representation for dynamic scenes and achieve significant acceleration over DyNeRF. HyperReel [1] proposes a network to predict the sample locations and represent dynamic scenes by key-frame-based TensoRF [9] representation to ensure high fidelity and real-time rendering without any custom CUDA kernels. K-Planes [17] and HexPlane [7] factorize the whole 4D field to the product of 2D planes. Though differently implemented, they both achieve high-fidelity results with intermediate model size (200MB) and training time (100 minutes) on DyNeRF dataset. TineuVox [16] leverages the deformation MLP in D-NeRF and implements a 3D voxel grid with multi-scale interpolation. It achieves high fidelity and fast convergence on D-NeRF dataset. MixVoxels [45] leverages the standard deviation across the time of all pixels to estimate a variation field to decompose the dynamic scene into static and dynamic voxels. Equipped with different representations for static and dynamic objects, MixVoxels achieves high-fidelity results and super-fast convergence (15 minutes) on DyNeRF dataset. Song et al. [43] train Instant-NGP [31] and TensoRF [9] offline and stream the models along the feature channels for dynamic scenes. Concurrent works are leveraging 4D Gaussian point clouds [54, 49] or 3D Gaussians with deformation networks [55] achieving real-time rendering and high fidelity in dynamic scenes.
@@ 3.1 Preliminaries
Let $I_{i}$ denote the $i^{th}$ input image corresponding to observation direction $d_{i}$, $\mathbf{V}=\{\nu_{i}~{}|~{}\nu_{i}=(I_{i},d_{i}),1\leq i\leq N_{view}\}$ is the set of the training views, where $i,N_{view}\in\mathbb{Z^{+}}$ and $\mathbf{D_{a}}$ is the readily-accessible data at one step. Previous methods [6, 35] for static scenes define the continual learning problem as learning a unified representation with a limited subset of views $V_{a}$ available at every optimization step, i.e. $\mathbf{D_{a}}=V_{a}\subsetneq V$. However, a more natural definition of continual NeRF is in the dynamic scenes, where the observations at different time stamps arrive continually. Denoting all the frame times as $\mathbf{T}=\{\tau_{i}~{}|~{}1\leq i\leq N_{time}\},i,N_{time}\in\mathbb{Z^{+}}$, we define the continual learning problem as synthesizing a video sequence when the model could only have access to a limited number of frames $\mathbf{T_{a}}\subsetneq\mathbf{T}$ from all views $\mathbf{V}$ at each step, i.e. $\mathbf{D_{a}}=\{\mathbf{V}\times\mathbf{T_{a}}\}$ where $\times$ is the Cartesian product of sets.
To simplify the notations, let the whole sequence of frames be evenly divided into $k$ chunks where a chunk $\mathbf{T_{k}}=\{\tau_{i}~{}|~{}\lfloor\frac{N_{time}}{T_{chunk}}\rfloor=k\}$ is a sequential subset of length $T_{chunk}$. We define the continual learning problem as learning a representation where there are only the frames in the current chunk available at each optimization step, i.e. $\mathbf{T_{a}}=\mathbf{T_{k}}$. Accordingly, $T_{chunk}=1$ denotes training the model frame by frame, $T_{chunk}=10$ denotes training the model with a buffer of 10 frames, $T_{chunk}=300$ denotes training the entire 300-frame sequence offline. The definition is close to real-world applications since caching all training frames in memory or storing them on the disk is limited in scalability for long videos.
@@ 3.2 Representation of Short Dynamic Intervals
As pointed out by [46], naive 4D hash grids are not suitable for high-fidelity dynamic view synthesis. To achieve fast reconstruction and slim model size simultaneously, we represent time and spatial features separately and concatenate them to alleviate feature confusion. The spatial hash features may be encoded along 3D voxels [31], 2D planes [17, 21], or both [38]. The time feature is encoded per frame following [28]. Following previous state-of-the-art methods [30, 31, 16], we apply a two-fold field function for geometry and color individually. The architecture of the field function is thus defined as:
@@
where $\Psi(\mathbf{x})$ denotes configurable spatial hash encodings, $f_{{\theta_{1}}}$ is a tiny MLP with trainable parameter ${\theta_{1}}$, $\Gamma(t)$ is the hash encoding for time axis alone, $\oplus$ is the concatenation of vectors. The MLP $f_{{\theta_{1}}}$ in Eq. 3 fuses features across the spatial domain and time domain and obtains the volume density $\sigma$. View-dependent colors $\mathbf{c}$ are regressed by another tiny MLP with parameter $\theta_{2}$ in Eq. 4. The spatial encoders $\Psi(x)$ follow a multi-resolution concatenated format, defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}~{}P_{m}(x,y,z)$ (5)
where $P_{m}(x,y,z)\in\mathbf{R^{F}}$ denotes the tri-linearly interpolated value from the nearby 8 grid vertices w.r.t. 3D coordinates $(x,y,z)$ at the $l^{th}$ resolution in Instant-NGP [31], $L$ is the number of the resolutions. $\Psi(x)$ may also be configured as the fusion of 2D and 3D features in MERF [38], defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}~{}P_{m}(x,y,z)+Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x)$ (6)
where $Q_{mij}(i,j)$ denotes the bi-linearly interpolated value from the 4 nearby vertices in the orthogonally projected planes w.r.t. axes $(i,j)$ at the $m^{th}$ resolution. Removing the $P_{m}(x,y,z)$ term leads to a pure plane representation like [17, 21].
3.3 CD-NGP
Unlike the existing replay-based methods for static scenarios [6, 35], we solve the continual learning problem of dynamic novel view synthesis via parameter isolation. The key insight is that the current state-of-the-art methods by tensor factorization [7, 17] or explicit voxels [45] are not accurate estimations of the object trajectory and inherent dynamics, but only simple regressions of the optical attributes in 3D space within a limited time. Therefore, the replay method is of limited use because the new transient processes are constantly arriving and the model of a limited size is not capable of reconstructing unlimited transient effects and thus suffers from catastrophic forgetting.
To address this issue, we propose our Continual Dynamic neural graphics primitives (CD-NGP) based on parameter isolation. We split the model into $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly. The $k^{th}$ branch contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$. We train the first branch (base branch) with a large spatial hash table to better reconstruct the scenes with initial movements and static objects. Then we reuse the spatial features from the base branch and train the subsequent branches (auxiliary branches) to compensate for residuals. We thus substitute the $\Psi(\mathbf{x})$ in Eq. 3 with the fused feature to obtain the volume density $\sigma$ at the $k^{th}$ model branch under the continual learning setting as:
@@ 3.3 CD-NGP
The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes. As pointed out by [45], most objects ($>90\%$) in the dynamic scenes are static. Therefore, the first chunk could be represented by a large hash table to reconstruct the scenes with initial movements and static objects, the subsequent chunks could be represented by small hash tables to compensate for errors, and the whole representation could be designed to leverage the fusion of features from the corresponding two branches to ensure scalability. To be specific, if we follow the default setting of Instant-NGP [31] and set the hash table in the first branch as $2^{P_{1}}=2^{19}$, the sizes of hash tables in subsequent branches could be configured as $2^{P_{2}}=2^{14}$, which consumes only $1/32$ of the parameters in the first branch. Since the spatial hash tables consume the most parameters in the model, and most chunks are encoded by the subsequent branches, the advantage of our asymmetric hash table structure is obvious: compared with the previous linear-streaming method [43], the total bandwidth consumption is reduced from $\mathcal{O}(\mathbf{K}\cdot 2^{19}LF)$ to $\mathcal{O}(N_{chunk}\cdot 2^{14}LF+2^{19}LF)$, where $\mathbf{K}$ is their bandwidth budget controlling the bitrates in streaming scenarios. If the bandwidth is extremely limited, the hash tables in our model could be arranged as Eq. 6 to reduce the bandwidth cost additionally.

### v3

@@ Abstract
We present CD-NGP, a fast and scalable representation for 3D reconstruction and novel view synthesis in dynamic scenes. Inspired by continual learning, our method first segments input videos into multiple chunks, followed by training the model chunk by chunk, and finally, fuses features of the first branch and subsequent branches. Experiments on the prevailing DyNeRF dataset demonstrate that our proposed novel representation reaches a great balance between memory consumption, model size, training speed, and rendering quality. Specifically, our method consumes $85\%$ less training memory ($<14$GB) than offline methods and requires significantly lower streaming bandwidth ($<0.4$MB/frame) than other online alternatives. Furthermore, we conduct experiments on a self-collected extra-long video sequences dataset, which demonstrates the superior scalability of our method over previous state-of-the-art approaches. A free copy of this paper and all supplemental materials are available at https://arxiv.org/abs/2409.05166.
@@
Continual learning, Dynamic novel view synthesis, Neural Radiance Field.
[…]
The demonstration of our continual learning pipeline for dynamic scenes. Please note the dimensions of representations for different branches remain the same while the complexities of hash tables are flexible and could be configured to ensure scalability. The spatial hash features could be encoded along voxels [31], planes [17] or both [38].
Introduction
@@
To address the dynamic reconstruction problem, a number of efforts have been made in the offline setting, requiring access to the entire dataset during the training stage. For example, the methods in [36, 28, 34, 33, 15, 29, 44] propose to use MLP-based representation and achieve high-quality results. There are also methods speed up the optimization by using tensor factorization ([40, 17, 7]) and voxel representation [16, 45]. The more recent continual-learning-based methods [27, 48] can train models frame by frame, and streamable representations [43] enable the model to be loaded on the fly at inference. However, the offline methods require huge host memory e.g., taking more than $100GB$ when reconstructing multi-view dynamic videos of 300 frames, and the continual learning-based methods are weak in balancing the convergence speed and model compactness. As a result, these methods are limited in scalability, making them unsuitable for long videos.
@@
In this paper, we propose continual dynamic neural graphics primitives (CD-NGP), a fast and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos. As illustrated in CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes, we evenly split the dynamic videos into multiple chunks based on the timestamp and conduct reconstruction with different model branches, where each branch includes MLP regressors, spatial encoding for the spatial coordinates $(x,y,z)$, and temporal encoding for the time stamp $t$. We encode the different chunks by hash tables with different complexities to ensure the model scalability and then continually train the model chunk by chunk to alleviate memory occupation. For each individual chunk, we locate the corresponding model branch, sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids [31], and represent spatial features by spatial hash grids. The spatial features of the first branch and the current branch are fused by element-wise sum as default or concatenation if applicable. The temporal features are encoded by the hash table for time stamp $t$ in the current branch only. Spatial and temporal features are concatenated along with viewing direction $d$ and input to MLP regressors for rendering and optimization.
We validate the effectiveness of our method through comprehensive experimental results on both the widely used public datasets and the dataset proposed by us. As illustrated in Fig. 1, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly, in which we use metrics on the flame salmon scene in the DyNeRF dataset for fair comparison. The perceptual consistency is measured by $1-$ LPIPS by AlexNet [57, 24], and the speed is the reciprocal of the training time. The compactness and RAM-Efficiency are mapped by $-log$ function from the model size and memory occupation respectively.
@@
Compared with prevailing offline baselines, our method consumes $85\%$ less host memory at the training phase, $40\%$ less model size at the inference phase, and delivers comparable results. Compared with online baselines, our method achieves compact model size, high quality, and fast convergence simultaneously. Contributions of this paper are three-fold as follows:
@@
  • We propose a novel method combining spatial and temporal hash encodings for dynamic scenes and yield comparable results to both offline and online methods.
@@
  • We propose a fast continual learning framework for dynamic scenes and reduce memory and storage occupation significantly.
@@
  • We provide a dataset of multi-view long video sequences ($>3$ minutes) with large motion and further demonstrate the scalability of our method on the proposed dataset.
@@ 1.1 Fast novel view synthesis in static scenes
The acceleration of NeRFs falls into two categories: training acceleration and rendering acceleration. Plenoxels [18] uses voxel grids to represent the scene and achieves magnitudes of acceleration over MLP backbones. TensoRF [9] decomposes the 3D scene as the product of tensors and achieves fast novel view synthesis. Plane-based methods [17, 21] project 3D points onto 3 orthogonal 2D planes to ensure sparsity and achieve fast convergence. Instant-NGP [31] leverages an occupancy grid and a multi-resolution hash encoding along with optimized tiny MLPs to enable fast novel view synthesis within a minute. The recent 3D Gaussian method [22] represents the scenes as a union of explicit Gaussian point clouds and achieves high fidelity, fast convergence, and real-time rendering on high-end hardware by eliminating the querying burden of spatial representations. KiloNeRF [37] distills the large MLP in NeRF [30] into tiny MLPs to accelerate rendering by thousands of times. Hedman et al. [20] propose the baking method (converting the MLP representation into explicit meshes) to accelerate the rendering process. They alpha-compose the cached features along the ray and use MLPs only once per ray to overcome the computation load of the per-point MLP querying in NeRF. BakedSDF [56] stores the view-dependent features into Gaussian spheres along the reconstructed surfaces, and achieves real-time rendering on ordinary laptops. MobileNeRF[10] represents scenes into meshes and enables real-time rendering on mobile devices. Light field methods [42, 8, 19] directly represent the scene as a set of rays to eliminate volume rendering and enable real-time applications on mobile devices.
@@ 1.2 Representations of dynamic scenes
There are mainly 3 variants of NeRF for dynamic scenes. The first class is pure MLP-based representation. D-NeRF [36] leverages a large deformation MLP to model the transition and the trajectories of the objects and extends NeRF to simulated dynamic scenes. DyNeRF [28] uses long latent vectors to encode temporal information along with large MLPs to achieve high-fidelity novel view synthesis in real dynamic scenes. Park et al. [34] propose an interpolation-based MLP representation for dynamic scenes and achieve significant acceleration over DyNeRF.
The second class is tensor factorization-based representation. HyperReel [1] proposes a network to predict the sample locations and represent dynamic scenes by key-frame-based TensoRF [9] representation to ensure high fidelity and real-time rendering without any custom CUDA kernels. K-Planes [17] and HexPlane [7] factorize the whole 4D field into the product of 2D planes. Though differently implemented, they both achieve high-fidelity results with intermediate model size (200MB) and training time (100 minutes) on the DyNeRF dataset. The third class is the explicit voxel-based representations. TineuVox [16] leverages the deformation MLP in D-NeRF and implements a 3D voxel grid with multi-scale interpolation. It achieves high fidelity and fast convergence on the D-NeRF dataset. MixVoxels [45] leverages the standard deviation across the time of all pixels to estimate a variation field to decompose the dynamic scene into static and dynamic voxels. Equipped with different representations for static and dynamic objects, MixVoxels achieves high-fidelity results and super-fast convergence (15 minutes) on DyNeRF dataset. Song et al. [43] train Instant-NGP [31] and TensoRF [9] offline and stream the models along the feature channels for dynamic scenes. Concurrent works are leveraging 4D Gaussian point clouds [55, 49] or 3D Gaussians with deformation networks [54] achieving real-time rendering and high fidelity in dynamic scenes.
@@ 2.1 Preliminaries
[…]
To simplify the notations, let the whole sequence of frames be evenly divided into $k$ chunks where a chunk $\mathbf{T_{k}}=\{\tau_{i}\leavevmode\nobreak\ |\leavevmode\nobreak\ \lfloor\frac{N_{time}}{T_{chunk}}\rfloor=k\}$ is a sequential subset of length $T_{chunk}$. We define the continual learning problem as learning a representation where only the frames in the current chunk are available at each optimization step, i.e. $\mathbf{T_{a}}=\mathbf{T_{k}}$. Accordingly, $T_{chunk}=1$ denotes training the model online in a frame-by-frame manner, $T_{chunk}=10$ denotes training the model online with a buffer of 10 frames. If the video contains 300 frames in total, the $T_{chunk}=300$ denotes training the entire 300-frame sequence offline. The definition is close to real-world applications since caching all training frames in memory or storing them on the disk is limited in scalability for long videos.
@@ 2.2 Representation of short dynamic intervals
As shown in [46], naive 4D hash grids for dynamic scenes are short in reconstruction quality. To achieve fast reconstruction and slim model size simultaneously, we represent time and spatial features separately and concatenate them to alleviate feature confusion. The spatial hash features may be encoded along 3D voxels [31], 2D planes [17, 21], or both [38]. The time feature is encoded by short hash tables in a per-frame manner. Following previous state-of-the-art methods [30, 31, 16], we apply a two-fold field function for geometry and color individually. The architecture of the field function is thus defined as:
@@
where $\mathbf{x}=(x,y,z)$ is the spatial coordinates, $\Psi(\mathbf{x})$ denotes configurable spatial hash encodings, $f_{{\theta_{1}}}$ is a tiny MLP with trainable parameter ${\theta_{1}}$, $\Gamma(t)$ is the hash encoding for time axis alone, and $\oplus$ is the concatenation of vectors. The MLP $f_{{\theta_{1}}}$ in Eq. 3 fuses features across the spatial domain and time domain and obtains the volume density $\sigma$. View-dependent colors $\mathbf{c}$ are regressed by another tiny MLP with parameter $\theta_{2}$ in Eq. 4. The spatial encoders $\Psi(x)$ follow a multi-resolution concatenated format, defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z)$ (5)
where the $P_{m}(x,y,z)\in\mathbf{R}^{F}$ denotes the tri-linearly interpolated value from the nearby 8 grid vertices w.r.t. 3D coordinates $(x,y,z)$ at the $l$-th resolution, $L$ is the number of the resolutions, and $F$ is the dimensions of features $P_{m}(x,y,z)$, which are similar to those in Instant-NGP [31].
2.3 Continual dynamic neural graphics primitives (CD-NGP)
Unlike the existing replay-based methods for static scenarios [6, 35], we solve the continual learning problem of dynamic novel view synthesis via parameter isolation. The current state-of-the-art methods by tensor factorization [7, 17] or explicit voxels [45] are regressions of the optical attributes in 3D space within a limited time instead of the object trajectory and inherent dynamics. Therefore, the replay method is of limited use in our defined scenario: Since new transient processes are constantly arriving, the model of a limited size is not capable of reconstructing unlimited transient effects. Thus, the replay-based method suffers from catastrophic forgetting.
To address this issue, we propose our continual dynamic neural graphics primitives (CD-NGP) based on parameter isolation. We split the model into $N_{chunk}$ branches to represent different chunks of the dynamic scene accordingly. The $k$-th branch contains spatial hash encoder $\Psi_{k}(\mathbf{x})$, temporal hash encoder $\Gamma_{k}(t)$, and MLPs $f_{\theta_{1k}},f_{\theta_{2k}}$. We train the first branch (base branch) with a large spatial hash table to better reconstruct the scenes with initial movements and static objects. Then we reuse the spatial features from the base branch and train the subsequent branches (auxiliary branches) to compensate for residuals. We thus substitute the $\Psi(\mathbf{x})$ in Eq. 3 with the fused feature to obtain the volume density $\sigma$ at the $k$-th model branch under the continual learning setting as:
@@ 2.3 Continual dynamic neural graphics primitives (CD-NGP)
The scalability of our method derives from the redundancy of dynamic scenes and the flexibility of hash table sizes. As pointed out by [45], most points ($>90\%$) in the dynamic scenes are static. We thus leverage the shared static points across time and represent the first chunk with a large hash table and the subsequent chunks with small hash tables to ensure scalability. To be specific, if we follow the default setting of Instant-NGP [31] and set the hash table in the first branch as $2^{P_{1}}$, the sizes of hash tables in subsequent branches could be configured as $2^{P_{2}}$. Please note that $P_{1}$ is significantly larger than $P_{2}$ to ensure the scalability of the model. When ($P_{1},P_{2})=(19,14)$, the hash table in the second branch consumes only $1/32$ of the parameters in the first branch. Since the spatial hash tables occupy the most parameters in the model, and most chunks are encoded by the subsequent branches, our assymetric hash table design enables continual learning of dynamic scenes and high scalability in streaming scenarios simultaneously. The total bandwidth consumption of our method is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, while the previous method [43] requires offline training and consumes $\mathcal{O}(\mathbf{K}\cdot 2^{P_{1}}LF)$ bandwidth, where $\mathbf{K}$ ranging from $0.5$ to $16$ is their bandwidth budget controlling the bitrates in streaming scenarios. If the bandwidth is extremely limited, the hash tables in our model could be arranged as a combination of 3D representations and 2D representations like [38] to reduce the bandwidth cost additionally.